# secReferences in the Online Appendix

### v2

@@ 6.2 NGO Success and Reputation
$\text{Success}_{i,n,y}$ $=\beta_{1}\text{Campaign on Same Topic}_{c(i,n,g),y}+\beta_{2}\text{Experience}_{g,y}$ (6)
$+\beta_{3}\text{Campaign on Same Topic}_{c(i,n,g),y}\cdot\text{Experience}_{g,y}$
@@
where $\text{Success}_{i,n,y}$ is defined as in (4)—namely, either a successful vote or a withdrawn proposal. On the right-hand side, $\text{Campaign on Same Topic}_{i,n,y}$ is one if there has been a campaign by NGO $g$ against firm $n$ on topic $i$ in the ten months preceding the AGM and zero otherwise, and $\text{Experience}_{g,y}$ indicates the number of campaigns NGO $g$ had launched in the past. We control for all confounding variations at the month, firm-by-year, NGO-by-year, and topic-by-year through fixed effects. Standard errors are clustered by firm.
@@
  • Note: This table reports the estimated coefficients from the regression of a dummy indicating whether an E&S shareholder proposal is withdrawn before voting or passes after voting (Success) on a dummy indicating the occurrence of an NGO campaign on the same topic as the proposal in the 10 months preceding the AGM (Campaign on Same Topic), a dummy indicating whether the NGO has launched a campaign on a firm’s AGM day in the past (AGM Experience (0/1)), and their interaction. In Column 2, we also consider the number of campaigns the NGO has launched in the past, which were not on the firm’s AGM day (Non-AGM Experience), and its interaction with Campaign on Same Topic. In Column 3 consider the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (AGM Experience), its square (Squared AGM Experience), and their interaction with Campaign on Same Topic. All regressions also include Firm x Year, NGO x Year, and Topic × Year fixed effects. Clustered (firm) standard errors are reported between parentheses.
Column (1) of Table 1 reports the estimated $\{\beta_{1},\beta_{2},\beta_{3}\}$. Campaigning on the same topic as a shareholder proposal only weakly influences the outcome of the proposal, and the $\beta_{1}$ coefficient is not significantly different from zero. Reassuringly, the number of past campaigns affects a proposal’s success only if the NGO campaigns on a topic being discussed at the AGM ($\beta_{2}=0$ but $\beta_{3}>0$): such campaigns increase a proposal’s success by 8% on average.
Column (2) unpacks Experienceg,y by considering the timing of an NGO’s past campaigns. We replace $\text{Experience}_{g,y}$ for the variable $\text{AGM Experience}_{g,y}$, which is 1 if NGO $g$ launched at least one campaign on the day of the AGM in past years and zero otherwise. The interaction coefficient is significant and economically important. While a campaign on the same proposal topic increases its success at the AGM by 8pp (26%), the effect increases by 14pp (45%) when that NGO has a history of targeting AGMs.
@@
Finally, as in signaling games where there are decreasing returns to signal one’s type if the agent has already signaled the type in the past, we estimate decreasing returns from launching campaigns on AGM dates. Column (4) replaces $\text{Experience}_{g,y}$ for the cumulative count of the campaigns when NGO $g$ targeted an AGM date ($\text{AGM Experience}_{g,y}$) and its squared value: the last two rows of Column (4) find that the coefficients of the interactions between those variables and Campaign on Same Topic imply a concave relationship between AGM experience and AGM success. The results are consistent with Figure 4, which shows that only the very first campaigns happen close to an AGM date.
@@
Targeting AGMs helps an NGO driving support for shareholder proposals in future periods, but display decreasing returns.
7 Discussion
@@
Our results demonstrate that NGOs strategically use the timing of their campaigns to influence firms’ decisions. In particular, as media coverage is higher during AGMs, NGOs are likely to wait for the AGM date to launch their campaigns against target firms’ practices or strategies (Result 1). By doing so, NGOs prompt reactions from key stakeholders, such as shareholders, who are more likely to file a related proposal at the following AGM, and consumers, as the likelihood of boycotts increases if NGOs campaign on the AGM date (Result 2). The goal of these campaigns is to alter a firm’s strategy by damaging its reputation and showing stakeholders’ disappointment (Bakos and Dellarocas, 2011; Kitzmueller and Shimshack, 2012).
@@
However, campaigning before the AGM offers direct advantages. Specifically, when a shareholder proposal has already been filed at a firm, targeting that firm on the same issue increases the likelihood that the firm accepts the proposal outright—potentially leading to its withdrawal from the AGM ballot—or, at the very least, raises shareholder support for the proposal by an average of 8% (Result 3). Since shareholder proposals generally receive low support, early NGO campaigns can play a role in disciplining managers (Levit and Malenko, 2011). Importantly, these analyses control for NGO-by-year fixed effects and topic-by-year fixed effects, accounting for any time-varying confounding factors at the NGO level—such as experience or reputation—or at the cause level—such as specific news events—thus ensuring that globally renown NGOs (e.g., WWF) do not solely drive these results. Also, we provide evidence suggesting that NGOs campaigns on AGMs do not target different firms compared to other campaign times in Appendix Table A2.
@@
Therefore, NGOs face a trade-off, which they navigate by adapting their strategies over time. NGOs with limited activism histories tend to campaign on the AGM date, while more established NGOs prefer targeting firms just before the AGM to directly influence shareholder votes (Result 4). This pattern is consistent with the idea of using campaigning time as a signaling device. One interpretation of this result is that, as NGOs gain more experience with campaigning—and potentially run multiple campaigns in a year—they accumulate exposure through standard news media and may already have a large social media following. As a result, there may be less pressure to campaign specifically on the AGM date. Another interpretation is that, as NGOs become more complex organizations over time, media exposure becomes less vital for influencing decision-making at target firms. Disentangling these interpretations suggest several potential avenues for future research, contributing to the broader conversation on civil society lobbying and activism (Bertrand et al., 2021, 2023).
The public may also expect mature NGOs not only to generate negative publicity but also to produce immediate results. Thus, there may be diminishing returns to reputation from consistently targeting AGMs. We show that this practice enables NGOs with a history of targeting AGMs to help related shareholder proposals pass. However, an excessively long history of targeting AGMs, which according to our estimates happens after 13 campaigns on average, decreases support for shareholder proposals (Result 5).
In conclusion, NGOs play a crucial role in increasing information and aligning corporate interests with societal needs, and their growth in recent years is well-documented (e.g., Egorov and Harstad, 2017; Daubanes and Rochet, 2019). However, our understanding of the incentives they face, their relationships with firms and stakeholders, and the strategies available to them remains limited due to the scant data available on NGOs and models to interpret their actions. This paper offers a detailed analysis of NGO campaigns to uncover these trade-offs and incentives by leveraging the exogeneity of the corporate calendar to NGO’s goals. Similar approaches have been successfully used to identify the preferences of other stakeholders and may help reframe the role of AGMs not just as voting venues but also as forums where firms engage with and negotiate alongside outside stakeholders in ways that are not often observed by researchers but can be inferred from the underlying system of incentives (Fioretti et al., 2023). We also acknowledge that NGOs not only campaign but also collaborate with corporations (Dahan et al., 2010). Future research could capitalize on similar exogenous variation to that employed in this paper to quantitatively study the incentives for collaboration.
By focusing on the timing of NGO campaigns, we demonstrate how this seemingly innocuous decision has significant implications for both targeted firms and the success of future initiatives. We show that, similar to for-profit organizations, NGOs are influenced by their reputation, which affects the timing of the information presented to stakeholders and, consequently, their responses. Thus, systematically studying NGOs’ objectives and how they evolve over time is essential for understanding whether and how their activism can help firms internalize firms’ externalities.
8 Conclusion
This paper examines the timing and motives behind NGO campaigns, revealing that these campaigns are not randomly timed but often strategically target firms’ Annual General Meetings. While this timing does not affect votes on environmental and social proposals, which are cast beforehand, it enhances the NGO’s visibility, leading to greater consumer reactions and increasing the likelihood of related shareholder proposals in the following year. Mature NGOs utilize the reputation similarly built over the years to influence current firm decisions by supporting relevant shareholder proposals. Thus, this paper contributes by empirically showing that an NGO’s objectives vary throughout its life, indicating that NGOs pursue different outcomes for targeted firms, their stakeholders, and society at large as they gain visibility, thereby fostering diverse discussions within companies and, potentially, in the broader public arena.
@@ Appendix B Omitted Figures
Figure B1: E&S shareholder proposals, and their success
@@
Note: This figure plots the estimated coefficients of the regression of a dummy indicating the occurrence of a shareholder proposal on month $t$ at firm $f$, relative to an NGO campaign targeting firm $f$ on the same topic, $k$ months ahead. The regression also includes month-by-year-by-industry-by-topic fixed effects. The regression is based on 1,227 NGO campaigns. Standard errors are clustered at the firm level.
@@ Appendix C Matching Proposals And NGO Campaigns
This section outlines the process of matching each shareholder proposal in the ISS dataset with a relevant NGO campaign topic.
Sigwatch assigns a“Generic Topic” to each campaign in their dataset. To streamline the analysis, we combine certain topics when they are difficult to distinguish or when one category has few campaigns and can reasonably fit within a broader category. Specifically, we merge “Labor Rights” with “Human & Labor Rights,” “Children” with ”Human & Labor Rights,” “CSR / ESG” with “CSR,” and “Biodiversity” and “Wildlife” with ”Environment.” For campaigns under the “Safety” topic, which includes only 12 entries, we examine each campaign’s specific purpose and classify it as either “Product Safety” or “Food Safety.”
Using these consolidated “Generic Topic Names”, we then categorize each shareholder proposal based on its short description. When a shareholder proposal is submitted, it includes a brief, one-sentence description, which is provided in the ISS dataset and serves as the foundation for topic classification. We start by assigning proposals to topics based on keywords in their descriptions and then manually classify any remaining proposals.
Table C1: Proposal Topics

### v3

@@ 6.2 NGO Success and Reputation
$\text{Success}_{i,n,y}$ $=\beta_{1}\,\text{Campaign on Same Cause}_{c(i,n,g),y}+\beta_{2}\,\text{Experience}_{g,y}$ (6)
$+\beta_{3}\,\text{Campaign on Same Cause}_{c(i,n,g),y}\cdot\text{Experience}_{g,y}$
@@
where “$\text{Success}_{i,n,y}$” is defined as in (4)—namely, either a successful vote or a withdrawn proposal. On the right-hand side, “$\text{Campaign on Same Cause}_{i,n,y}$” is one if there has been a campaign by NGO $g$ against firm $n$ on cause $i$ in the ten months preceding the AGM and zero otherwise, and “$\text{Experience}_{g,y}$” indicates the number of campaigns NGO $g$ had launched in the past. We control for all confounding variations at the month, firm-by-year, NGO-by-year, and cause-by-year through fixed effects. Standard errors are clustered by firm.
@@
  • Note: This table reports the estimated coefficients from (6) in Column 1. In Column 1, Experience is capped at 5 to avoid outliers–as 85% of NGOs never reach that level of experience. In Column 2, we also consider the number of campaigns the NGO has launched in the past, which were not on the firm’s AGM day (Non-AGM Experience), and its interaction with Campaign on Same Topic. In Column 3 consider the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (AGM Experience), its square (Squared AGM Experience), and their interaction with Campaign on Same Topic. All regressions also include firm-by-year, NGO-by-year, topic-by-year, and month fixed effects. Clustered (firm) standard errors are reported in parentheses.
Column (1) of Table 1 reports the estimated $\{\beta_{1},\beta_{2},\beta_{3}\}$. Campaigning on the same cause as a shareholder proposal only weakly influences the outcome of the proposal, and the $\beta_{1}$ coefficient is not significantly different from zero. Reassuringly, the number of past campaigns affects a proposal’s success only if the NGO campaigns on a cause being discussed at the AGM ($\beta_{2}=0$ but $\beta_{3}>0$): such campaigns increase a proposal’s success by 8% on average.
Column (2) unpacks Experienceg,y by considering the timing of an NGO’s past campaigns. We replace “$\text{Experience}_{g,y}$” for the variable “$\text{AGM Experience}_{g,y}$,” which is 1 if NGO $g$ launched at least one campaign on the day of the AGM in past years and zero otherwise. The interaction coefficient is significant and economically important. While a campaign on the same cause as the shareholder proposal increases its success at the AGM by 8pp (26%), the effect increases by 14pp (45%) when that NGO has a history of targeting AGMs.
@@
Finally, as in signaling games where there are decreasing returns to signal one’s type if the agent has already signaled the type in the past, we estimate decreasing returns from launching campaigns on AGM dates. Column (4) replaces “$\text{Experience}_{g,y}$” for the cumulative count of the campaigns when NGO $g$ targeted an AGM date (“$\text{AGM Experience}_{g,y}$”) and its squared value: the last two rows of Column (4) find that the coefficients of the interactions between those variables and “$\text{Campaign on Same Cause}_{i,n,g,y}$” imply a concave relationship between past AGM-targeting and proposals’ success. Therefore, the support of NGO’s to a shareholder proposal display diminishing returns.
This finding supports the results in Appendix Figure B4, discussed in Section 4, which demonstrate that AGM-targeting increases the probability that target receives negative media coverage only if the NGO has little campaign experience (i.e., having run less than four campaigns). Later campaigns do not generate additional awareness on AGM dates compared to other dates. In fact, “mature NGOs” seem to be better off campaigning one or two months earlier, indicating diminishing returns from campaigning on AGM dates, even in terms of visibility.
We conclude that,
@@
Targeting AGMs helps an NGO driving support for shareholder proposals in future periods, but displays decreasing returns.
7 Discussion and Conclusion
By analyzing the timing of NGO campaigns, we demonstrate that this seemingly minor decision has substantial consequences for both targeted firms and the effectiveness of future initiatives. Furthermore, we demonstrate that NGOs’ tactics evolve over time as they seek to either indirectly pressure firms (“exposure”) or directly shape their operations (“influence”). Thus, systematically analyzing NGOs’ behavior can shed new light on how activism helps firms internalize externalities.
@@ 7 Discussion and Conclusion
We first show that NGOs tend to launch their campaigns on the AGM date, as media coverage is higher during AGMs (Result 1). By doing so, NGOs elicit reactions from key stakeholders, such as shareholders and consumers (Result 2). The immediate goal of these campaigns is to influence a firm’s strategy by damaging its reputation and voicing stakeholder dissatisfaction.
@@
However, campaigning before the AGM offers direct advantages. Specifically, when a shareholder proposal has already been filed at a firm, targeting that firm on the same issue increases the likelihood that the firm accepts the proposal outright—potentially leading to its withdrawal from the AGM ballot—or, at the very least, raises shareholder support for the proposal by an average of 8% (Result 3). Since shareholder proposals generally receive low support, early NGO campaigns can discipline managers (Levit and Malenko, 2011). Importantly, all these analyses control for NGO-by-year fixed effects and cause-by-year fixed effects, accounting for any time-varying confounding factors at the NGO level—such as experience or reputation—or at the cause level—such as specific news events—thus ensuring that globally renown NGOs (e.g., PETA) do not drive these results. Also, we provide evidence suggesting that NGO campaigns on AGMs do not target different firms compared to campaigns at other times, which helps mitigate selection issues (Appendix Table A2).
@@
Therefore, NGOs face a trade-off, which they navigate by adapting their strategies over time. NGOs with limited activism histories tend to campaign on the AGM date, whereas more established NGOs prefer targeting firms just before the AGM to directly influence shareholder votes (Result 4). One interpretation is that as NGOs gain more experience with campaigning, they accumulate media exposure, reducing the pressure to campaign specifically on the AGM date.
Therefore, repeatedly targeting AGMs may yield diminishing returns (Result 5). First, an NGO may have already signaled its commitment to a cause, making subsequent signals weaker and reducing the effectiveness of AGM-date campaigning as a signaling device. This interpretation aligns with the view that AGM-date campaigns enhance an NGO’s “reputation.” For instance, following ”reputation-for-toughness” models, the NGO may accept an opportunity cost—delaying support for a shareholder proposal—to launch a campaign at a time when it imposes the greatest costs on a firm, thereby building a reputation that can be leveraged with other targets in the future. [^13] In these models, incumbent firms may incur short-term losses by setting predatory prices upon entry of a rival to deter future entrants (Kreps and Wilson, 1982; Milgrom and Roberts, 1982).
Second, stakeholders may come to expect more constructive approaches and proposals from established NGOs, leading NGOs to support earlier shareholder proposals or other unobserved tactics. Indeed, we find that mature NGOs—those with multiple prior campaigns in the sample—no longer inflict additional reputation losses by campaigning on the AGM date, whereas they did so in their first campaigns (compare the blue dots and red triangles in Appendix Figure B4). In fact, experienced NGOs generate the same or even more negative media coverage for a firm by launching campaigns one or two months before the AGM.
Other interpretations are unlikely to be relevant unless they covary with the number of AGM-date campaigns within a given year, NGO, firm, and cause. For instance, organizational aging (e.g., older NGOs becoming larger and more complex, with multiple ways to influence firms) is accounted for through fixed effects. These interpretations open avenues for future research, contributing to the broader discussion on civil society lobbying and activism (Bertrand et al., 2021, 2023).
In conclusion, NGOs influence corporate behavior, yet their incentives and strategies remain understudied. This paper analyzes NGO campaigns, leveraging the exogeneity of the corporate calendar to uncover their trade-offs. Similar approaches have identified stakeholder preferences, showing that AGMs are more than voting venues—they also serve as forums for firm-stakeholder negotiations, often unseen by researchers but inferred from underlying incentives (Fioretti et al., 2023).
We also acknowledge that NGOs not only campaign but also collaborate with corporations (Dahan et al., 2010). Future research could capitalize on similar exogenous variation to quantitatively study the incentives for collaboration.
@@ Appendix B Omitted Figures
Figure B1: Count of E&S shareholder proposals and their success
@@
Figure B4: Media exposure of NGO campaigns around an AGM
(a) Main specification
(b) Controlling also for current E&S news
Note: The figure presents estimates from (1). The dependent variable is an indicator that a firm receives at least one negative E&S news in a week. Panel (b) also controls for the current count of E&S news to account for firm-week specific events that increase news coverage (not necessarily negative coverage). Standard errors are clustered at the firm level.
Figure B5: Examples of positive and negative E&S news
(a) Example of a positive E or S news
(b) Example of a negative E or S news
Notes: This figure displays examples of a positive (top) and negative (bottom) E&S news. News data from RavenPack.
Figure B6: Media sentiment of NGO campaigns around an AGM
Note:The figure presents estimates from (1). The dependent variable is the average sentiment of E&S news received by firm in a week. Standard errors are clustered at the firm level.
@@
Note: This figure plots the estimated coefficients of the regression of a dummy indicating the occurrence of a shareholder proposal on month $t$ at firm $f$, relative to an NGO campaign targeting firm $f$ on the same cause, $k$ months ahead. The regression also includes month-by-year-by-industry-by-topic fixed effects. The regression is based on 1,227 NGO campaigns. Standard errors are clustered at the firm level.
@@
Figure B9: Campaign Timing and E&S Score Improvement
Note: This figure plots the estimated coefficients of the following regression:
$\Delta_{4}\text{E\&S Score}_{n,q}=\sum_{l=-6}^{5}\text{Campaign}_{n,q,l}+\textbf{X}_{n,q}\gamma^{\prime}+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$
where $\Delta_{4}\text{E\&S Score}_{i,q}$ is the change in firm $n$’s average of its Environmental and Social scores between quarter $q$ and quarter $q+4$, and $\sum_{l=-6}^{6}\text{Campaign}_{n,q,l}$ is a set of dummy variables indicating the occurrence of an NGO campaign against firm $n$ in quarter $q$, being $l$ months away from the closest AGM. The regression also controls for firm size (logged total assets), current average of E and S scores, firm fixed effects ($\alpha_{n}$), and industry-by-quarter fixed effects ($\alpha_{s(n),q}$). ESG scores are from MSCI, NGO campaigns data from Sigwatch, firm characteristics from Compustat, and AGM dates from ISS. The regression has 47,082 observations. Standard errors are clustered at the firm level.
@@ Appendix C Matching Proposals And NGO Campaigns
This section outlines the process of matching each shareholder proposal in the ISS dataset with a relevant NGO campaign cause.
Sigwatch assigns a“Generic Topic” to each campaign in their dataset. To streamline the analysis, we combine certain causes when they are difficult to distinguish or when one category has few campaigns and can reasonably fit within a broader category. Specifically, we merge “Labor Rights” with “Human & Labor Rights,” “Children” with ”Human & Labor Rights,” “CSR / ESG” with “CSR,” and “Biodiversity” and “Wildlife” with ”Environment.” For campaigns under the “Safety” cause, which includes only 12 entries, we examine each campaign’s specific purpose and classify it as either “Product Safety” or “Food Safety.”
Using these consolidated “Generic Cause Names”, we then categorize each shareholder proposal based on its short description. When a shareholder proposal is submitted, it includes a brief, one-sentence description, which is provided in the ISS dataset and serves as the foundation for cause classification. We start by assigning proposals to causes based on keywords in their descriptions and then manually classify any remaining proposals.
Table C1: Proposal Causes